John Smith is about to retire, and he wants to buy an annuity that will provide him with $50,000 of income per year for 20 years, beginning a year from today. The going rate on such annuities is 10%. How much would it cost him to buy such an annuity today? (Sample Answer: $350,789.25)
Answer #1

Cost of annuity today is $ 425,678.19

Present value of annual cash flow = Annual cash flow * Present value of annuity of 1
= $       50,000.00 * 8.513564
= $   4,25,678.19
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.10)^-20)/0.10 i = 10%
= 8.51356372 n = 20